OpenStudy (anonymous):

6. If x − y/x + y = 7 (x and -y aren't equal), what is the ratio of y to x? (A) −4/3 (B) 4/3 (C) −3/4 (D) 3/4

OpenStudy (anonymous):

is it (x-y)(x+y)=7?

OpenStudy (anonymous):

I'm not sure the answer is C but I don't know how to solve it

Vocaloid (vocaloid):

lets start over (x-y)/(x+y) = 7

Vocaloid (vocaloid):

multiply both sides by (x+y)

OpenStudy (anonymous):

okay

OpenStudy (welshfella):

is the question x - y ---- = 7 ? x + y

OpenStudy (anonymous):

i think so @welshfella . @taengoo is what @welshfella wrote the exact question you have?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

after u do what @Vocaloid wrote u to do, you should get x-y=7(x+y) simplify this expression first

OpenStudy (welshfella):

vocaloids suggestion looks good

OpenStudy (anonymous):

x-y=7x+7y

OpenStudy (campbell_st):

mu;tiply both sides of the equation by x + y and you get x - y = 7(x + y) now distribute and collect like terms

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

yes. so now subtract x from both sides and 7y from both sides. x - y - x - 7y = 7x + 7y - x - 7y simplify this one

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

y=-3/4x

OpenStudy (welshfella):

right

OpenStudy (anonymous):

yes. that's right.

OpenStudy (anonymous):

Can you help me on another one?

OpenStudy (anonymous):

of course :)

OpenStudy (welshfella):

please re-post another question